The Smith Seed Company grows heirloom tomatoes and sells their seeds. The heirloom tomato plants are preferred by many growers for their superior flavor. At the end of the most recent year the firm had current assets of $49,800 net fixed assets of $250,100, current liabilities of $31,100, and long-term debt of $101,100.

a. Calculate Smith’s stockholders' equity

b. What is the firm's net working capital?

c. If Smith's current liabilities consist of $21,400 in accounts payable and $9,700 in short-term debt (notes payable), what is the firm's net working capital?
Answer #1

(a)Stockholder equity =Total assets -Total liabilities

=Current assets + net fixed assets-current liabilities -long term debt

=49800+250100-31100-101100

=$167700

(b)Net working capital =Current assets -Current liabilities

=49800-31100

=$18700

(C) given that current liabilities consists of accounts payable and notes payable and both these items are included in current liabilities while calculating Net working capital .Therefore,

Net working capital=current assets - current liabilities

=Current assets -(accounts payable+notes payable)

=49800-(21400+9700)

=$18700